Find the present value of $60,000 due in 4 yr at the given rate of interest. (Round your answer to the nearest cent.)
12%/year compounded quarterly

Answer by rfer(16322)

P=A/(1+.12/4)^4*4
P=60000/(1.03)^16
P=60000/(1.6047)
P=$37390.17